OpenStudy (anonymous):

Given the equation y − 4 = 3/4(x + 8) in point-slope form, identify the equation of the same line in standard form. −3/4x + y = 10 3x − 4y = −40 y = 3/4x + 12 y = 3/4x + 10

OpenStudy (anonymous):

y-4=3/4x + 3(8)/4 y-4 = 3/4 x + 6 y=3/4x + 6 +4 y=3/4 x + 10

OpenStudy (anonymous):

y-4=3/4x + 3(8)/4 y-4 = 3/4 x + 6 y=3/4x + 6 +4 y=3/4 x + 10